OpenStudy (anonymous):

if f^-1(x) is the inverse of f(x)=2e^-x, then f^-1(x)=

OpenStudy (anonymous):

f(x) = 2e^-x y = 2e^-x (rewrite f(x) as y x = 2e^-y (interchange x and y) ln(x) = ln(2e^-y) (take ln both sides to solve for y) ln(x) = ln2 + ln(e^-y) ln(x) = ln2 -y y = ln2 - lnx y = ln(2/x) so f inverse is ln(2/x)